# Networked Communication for Decentralised Cooperative Agents in Mean-Field Control

### v1

@@ 5.1 Experimental setup
Figure 1: Standard algorithm settings with $C_{e}=C_{r}=C_{p}=1$. In almost all games the centralised (blue) and independent (orange) agents are significantly outperformed by networked agents of all broadcast radii.
5.2 Results and discussion
We present results in Fig. 1 for our standard experimental settings involving 500 agents each with their own Q-network. When networked agents communicate, they have only a single communication round. See Appx. E.3 for additional experiments with more communication rounds, a study of robustness to failures in the communication network, and ablation studies for our various sub-routines. The ablation studies of Algs. 1 (estimating global average reward) and 3 (estimating global empirical mean field) suggest that in our experimental settings the policy communication scheme (Alg. 4) is the dominant factor in the better performance of networked populations over the other architectures.
@@ 5.2 Results and discussion
Fig. 1 shows that in all of our games, networked populations of all broadcast radii significantly outperform independent (orange) agents, which hardly appear to increase their returns, if at all. Networked populations of all broadcast radii also significantly outperform the centralised (blue) agents in all but the two coordination games, where only networked agents of the smaller radii (green, 0.2; red, 0.4; purple, 0.6) underperform them. Indeed, in the anti-coordination games the centralised populations perform similarly to purely independent ones in hardly appearing to increase their returns, performing even worse than independent ones in the ‘shape formation’ game. The centralised populations also have markedly higher variance than networked ones in several games (‘target selection’, ‘disperse’, ‘beach bar’). This reflects our theoretical analysis in Sec. 4 that the centralised learner pushes an arbitrary updated policy to the whole population regardless of its quality, leading to large fluctuations in performance, whereas our communication scheme biases networked populations towards better performing updates.
@@
In the four anti-coordination games, and most notably in the ‘target coverage’ game, networked agents of smaller broadcast radii generally outperform those of larger radii, i.e. the ordering is reversed from that of the coordination games. This reflects the fact that our strong theoretical Ass. 8 (namely that an increase in the base return function must outweigh a decrease in the population’s policy diversity in anti-coordination games) only applies to a certain extent in practice, as explained in the following.

### v2

@@ 6.3 Results and discussion
Figure 2: Standard settings with $C_{e}=C_{r}=C_{p}=1$. In all games networked agents of all broadcast radii significantly outperform the independent (orange) populations, and in most games they also outperform the central-agent (blue) populations, reflecting our theoretical results. The central-agent populations also have markedly higher variance than networked ones in several games, since the central learner pushes an arbitrary updated policy to the whole population regardless of its quality, leading to large fluctuations in performance, whereas our communication scheme biases networked populations towards better performing updates.
Fig. 2 gives results for our standard experimental settings involving 500 agents, each with their own Q-network. When networked agents communicate, they have only a single communication round. Fig. 2 shows that in all of our games, networked populations of all broadcast radii significantly outperform independent (orange) agents, which hardly appear to increase their returns, if at all. Networked populations of all broadcast radii also significantly outperform the central-agent (blue) populations in all but the two coordination games, where only networked agents of the smaller radii (green, 0.2; red, 0.4; purple, 0.6) underperform them (due to these less connected populations being more likely to experience violations of Assumption 5.7 on policy consensus, which is a disadvantage in scenarios where alignment is beneficial). Indeed, in the anti-coordination games the central-agent populations perform similarly to purely independent ones in hardly appearing to increase their returns, performing even worse than independent agents in the ‘shape formation’ game. The central-agent populations also have markedly higher variance than networked ones in several games (‘target selection’, ‘disperse’, ‘beach bar’). This reflects our theoretical analysis in Sec. 5 that the central learner pushes an arbitrary updated policy to the whole population regardless of its quality, leading to large fluctuations in performance, whereas our communication scheme biases networked populations towards better performing updates.
@@
In the ‘target coverage’ game, and sometimes the other anti-coordination games to a lesser extent, networked agents of smaller broadcast radii appear to outperform those of larger radii, i.e. the ordering is reversed from that of the coordination games, albeit not necessarily significantly so. This reflects the point up to which our Assumption 5.14 (increase in base return outweighs decrease in diversity in anti-coordination games, and base return is not yet maximised), holds in practice, which we discuss in the following.
The first part of Assumption 5.14 strictly holds throughout the ‘disperse’, ‘target coverage’ and ‘shape formation’ anti-coordination games: agents get no reward for diversity unless they are stationary (and also unless they are in one of the correct locations in the latter two cases). This means that any increase in base return (likelihood of being stationary or in the right location) achieved by policy adoption does indeed outweigh the loss of diversity. The first part of Assumption 5.14 mostly holds in the ‘beach bar’ game, apart from in a small window for agents that are stationary close to the bar target, with the window defined by the size of the population and hence the potential magnitude of the $\log(\hat{\mu}_{t}(s^{i}_{t})$ term in the reward function. Inside this window, increasing base return by moving even closer to the target, at the cost of being in a more crowded area, would not necessarily be beneficial. Regardless, in all these games the networked populations of all broadcast radii significantly outperform the independent agents, which do not appear to be able to learn at all without the helpful bias towards policies with better base returns enabled by the communication scheme.
However, among these networked populations, the base return quickly reaches its capacity, i.e. agents learn to be primarily stationary in one of the right locations, such that the second part of Assumption 5.14 no longer holds. This is not an issue when comparing with the independent populations, which have not maximised their base returns and therefore perform worse, but it does give rise to the reverse ordering of returns which we see among networked populations of different radii and hence connectivities. Once base return is maximised, policies that are estimated to receive higher returns may be less aligned with other policies than those other policies are with each other (at least regarding the strategically relevant parts of policies which are rewarded for greater diversity, e.g. these policies visit the less congested locations), or they simply visited the less congested locations by chance during the finite evaluation steps. Either way, more adoption of policies now becomes a disadvantage, since it reduces diversity without an additional positive impact on base return. Therefore architectures that give less communication now perform better by preserving diversity. Populations with lower broadcast radii usually have less connected networks, especially if sub-populations become isolated from each other, which is more likely in our ‘target coverage’ game than the others since the target locations are as far apart as possible from each other. Therefore these populations have less communication than those with larger broadcast radii and so may perform better, even while all networked populations outperform the independent agents that have not maximised their base returns.